OpenStudy (anonymous):

it is posible to exist the three diferent types of asymptotes in a single function?

OpenStudy (anonymous):

nope.

OpenStudy (anonymous):

Actually, hells yes.

OpenStudy (anonymous):

do you know why?

OpenStudy (anonymous):

I'll give you an example. Just a moment.

OpenStudy (anonymous):

How about this function: for x < 5, y=1/x for x >=5, y = x+1/x

OpenStudy (anonymous):

Blam. Horizontal asymptote to the left. Vetical asymptote at 0. Oblique asymptote to the right.

OpenStudy (anonymous):

Questions?

OpenStudy (anonymous):

thank you, that really helped

OpenStudy (anonymous):

My pleasure =)